OpenStudy (anonymous):

The numbers 20,21 and c are a Pythagorean triple . what is the value of c ? A. 22 B.41 C.29 D.25

OpenStudy (anonymous):

I will give you a medal :)

OpenStudy (kc_kennylau):

For a Pythagorean triple (x,y,z), \(x^2+y^2=z^2\)

OpenStudy (anonymous):

\[20^2+21^2=c^2~~~~~c=?\]

OpenStudy (anonymous):

Im so confused ....

OpenStudy (anonymous):

The formula for a Pythagorean triple is \[a^2+b^2=c^2\] look at the angle below...  where a and b are the legs of the triangle, and c is the hypotenuse. Every such a triangle must fit \[a^2+b^2=c^2\]

OpenStudy (anonymous):

In your case, you are told that 20, 21 and c is a Pythagorean triple, so it therefore must fit \[20^2+21^2=c^2\] correct ?

OpenStudy (anonymous):

Ok ? But theres numbers for the answers .

OpenStudy (anonymous):

yes, can you solve for c, in the equation below? \[20^2+21^2=c^2\]

OpenStudy (anonymous):

Ummmmm ... Im bad at math ....

OpenStudy (anonymous):

you are given  (I am not going to go off the way from assuming that C is the hypotenuse, because doing that, you won;t get any of the options) ANYWAY.... So, you will set the equation, \[20^2+21^2=c^2~~~~~~~~Correct?\]

OpenStudy (anonymous):

I am going to provide the steps to do these equations, and you will love it okay? Prepare....!

OpenStudy (anonymous):

1) take square root of each side of the equation ( 20^2+21^2=c^2 ). 2) solve (i.e. calculate) what you get inside the square root on the left side. 3) what is the square rot of the result that you get on the left side?